# On decoding extended Han-Zhang codes

### v1

@@ III-A The existence of $\ell$-ECPs of extended Han-Zhang codes
    If $2\nmid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\}~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty).$
@@
    If $2\mid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1).$
@@
1) Since $\mathcal{S}$} contains a $k$-zero-sum subset, it follows from Lemma 4.2) that $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1$. From $2\nmid(n-k)$ and $3\leq k\leq n-2$, we get that $3\leq k\leq n-3$ and $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k-1}{2}\leq\frac{n-4}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell-1,2\ell+2]_{q}$. In this case, we derive from Lemma 3 that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ has a generator matrix of the form
@@
Take $\mathcal{A}=\{(c_{1},\ldots,c_{n},0):~(c_{1},\ldots,c_{n})\in{\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1})\}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\}$ and $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)$. According to (1), we deduce that $\mathcal{A}$ and $\mathcal{B}$ have generator matrices of the forms
$G_{\mathcal{A}}=\begin{pmatrix}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a^{\ell}_{1}&\ldots&a^{\ell}_{n}&0\end{pmatrix}~{\rm and}~G_{\mathcal{B}}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-1}&1\\ \end{pmatrix}$,
@@
$\begin{split}d(\mathcal{B}^{\perp})&=d({\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)^{\perp})=\ell+1>\ell,\\ \dim(\mathcal{A})&=\ell+1>\ell,~{\rm and}\\ d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))&=n+1-(\ell+1)+2\ell+2=n+\ell+2>n+1.\end{split}$
@@
$(\mathcal{A},\mathcal{B})=\left(\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\},{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof of 1).
@@
$d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n+1-(\ell+1)+1+2\ell+1=n+\ell+2>n+1.$
@@
$(\mathcal{A},\mathcal{B})=\left({\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ by Definition 5. This completes the proof of 2). ∎
@@
Note that we suppose $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$ for a more convenient representation in Theorems 13 and 14. Combining with Lemma 4, we can further deduce the explicit values of $\ell$ as listed in Table I, where the notation is the same with that used in Theorems 13 and 14.
TABLE I: $\ell$-ECPs $(\mathcal{A},\mathcal{B})$ of extended Han-Zhang codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$
@@
III-B An explicit decoding algorithm of extended Han-Zhang codes based on their $\ell$-ECPs
In this subsection, we provide an explicit decoding algorithm for extended Han-Zhang codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on their $\ell$-ECPs. First of all, we recall the error-correcting mechanism of $\ell$-ECPs for a given $n,k]_{q}$ linear code $\mathcal{C}$} (see also [16, 8]). Let $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of $\mathcal{C}$} and let ${\bf y}={\bf c}+{\bf e}\in\mathbb{F}_{q}^{n}$ be a received vector with ${\bf c}\in{\mathcal{C}}$ and ${{\rm{wt}}}({\bf e})\leq\ell$. Note that based on conditions $\rm i)$$\rm iii)$} in Definition 5, Duursma and Kötter showed in [10, Theorem 1] that there exists a nonzero vector ${\bf a}\in\mathcal{A}$ such that
@@ III-B An explicit decoding algorithm of extended Han-Zhang codes based on their $\ell$-ECPs
has a unique solution, where $H_{{\mathcal{C}}}$ is a parity-check matrix of $\mathcal{C}$}. It then follows that ${\bf x}={\bf e}$ and $x_{i}=0$ for all $i\notin z({\bf a})$ if we write ${\bf x}=(x_{1},\ldots,x_{n})$. As a result, the $\ell$-ECP $(\mathcal{A},\mathcal{B})$ has decoded the received vector $\bf y$} to the codeword ${\bf c}={\bf y}-{\bf x}$.
Based on the above discussion, we conclude that to decode a linear code $\mathcal{C}$} using its $\ell$-ECP $(\mathcal{A},\mathcal{B})$, one should first determine explicit representations of the code pair $\mathcal{A}$ and $\mathcal{B}$, and then solve the system of equations (17) and (18). By combining Theorems 13 and 14, we can now present a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on its $\ell$-ECP as follows.
@@
If $\mathcal{S}$} is $k$-zero-sum free and $2\nmid(n-k)$, then an extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ can be decoded by Algorithm 1 in $O(n^{3})$ time. Moreover, the decoding algorithms with the same time complexity for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ with respect to the other three cases discussed in Theorems 13 and 14 can be similarly derived by modifying Lines 8-14 of Algorithm 1 based on Table I accordingly.
@@
The decoding correctness is straightforward by combining Theorems 13 and 14 as well as the discussions above. By [36, Theorem 2.14], the time complexity is $O(n^{3})$. More details can be found in Algorithm 1. ∎
@@ 11    ${{\mathbf{v}}}^{\prime}\leftarrow(v_{1}(a_{1}-\gamma)^{k-1},\ldots,v_{n}(a_{n}-\gamma)^{k-1},1)$
12    ${{\mathbf{u}}}^{\prime}\leftarrow(u_{1}^{\prime},\ldots,u_{n+1}^{\prime})$ with $u_{i}^{\prime}$ given in (13)
@@ 26          ${\bf c}\leftarrow{\bf y}-{\bf x}_{0}$
Algorithm 1 An algorithm for decoding ${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)$ when $\mathcal{S}$} is $k$-zero-sum free and $2\nmid(n-k)$
@@
Let $q=17$, ${\mathcal{S}}=\{1,3,5,7,10,12,14,16\}\subseteq\mathbb{F}_{17}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1)\in(\mathbb{F}_{17}^{*})^{8}$. Consider the extended Han-Zhang code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ with generator matrix
@@
Suppose that the received vector is ${\bf y}=(4,6,1,14,5,7,12,15,2)\in\mathbb{F}_{17}^{9}$ and we have
@@
  • Note that $\mathcal{S}$} is $3$-zero-sum free. It then follows from Lemmas 3, 4.1), and 5 that the extended Han-Zhang code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $9,3,7]_{17}$ non-GRS MDS code with parity-check matrix
@@
    $G_{\mathcal{A}}=\begin{pmatrix}1&1&1&1&1&1&1&1&1\\ 16&1&6&7&15&12&10&11&0\\ 1&1&2&15&4&8&15&2&0\\ \end{pmatrix}~\mbox{and}~G_{\mathcal{B}}=\begin{pmatrix}5&14&13&3&12&13&5&3&3\\ 12&14&10&4&10&3&16&16&0\end{pmatrix}$.
@@
    As a result, we get the solution ${\bf x}_{0}=\{0,16,0,0,0,6,0,0,0\}$ and ${{\rm{wt}}}({\bf x}_{0})=2$.
  • Finally, according to Lines 24 and 25 of Algorithm 1, the received vector $\bf y$} is then decoded to the codeword
@@
Let $q=16$ and let $\omega$} be a primitive element of $\mathbb{F}_{16}$ such that ${\omega}^{4}+{\omega}+1=0$, ${\mathcal{S}}=\{1,{\omega},{\omega}^{2},{\omega}^{3},{\omega}^{4},{\omega}^{6},{\omega}^{7},{\omega}^{8},{\omega}^{9},{\omega}^{11},\\ {\omega}^{12},{\omega}^{13},{\omega}^{14}\}\subseteq\mathbb{F}_{16}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1,1,1,1,1,1)\in(\mathbb{F}_{16}^{*})^{13}$. Consider the extended Han-Zhang code ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$ with generator matrix
$G_{7}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{3}&{\omega}^{4}&{\omega}^{6}&{\omega}^{7}&{\omega}^{8}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{13}&{\omega}^{14}&0\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{6}&{\omega}^{8}&{\omega}^{12}&{\omega}^{14}&{\omega}&{\omega}^{3}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{13}&0\\ 1&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&0\\ 1&{\omega}^{4}&{\omega}^{8}&{\omega}^{12}&{\omega}&{\omega}^{9}&{\omega}^{13}&{\omega}^{2}&{\omega}^{6}&{\omega}^{14}&{\omega}^{3}&{\omega}^{7}&{\omega}^{11}&0\\ 1&{\omega}^{5}&{\omega}^{10}&1&{\omega}^{5}&1&{\omega}^{5}&{\omega}^{10}&1&{\omega}^{10}&1&{\omega}^{5}&{\omega}^{10}&0\\ 1&{\omega}^{7}&{\omega}^{14}&{\omega}^{6}&{\omega}^{13}&{\omega}^{12}&{\omega}^{4}&{\omega}^{11}&{\omega}^{3}&{\omega}^{2}&{\omega}^{9}&{\omega}&{\omega}^{8}&1\\ \end{array}\right).$
@@
${\mathrm{diag}}({\bf y})={\mathrm{diag}}({\omega}^{12},{\omega},{\omega}^{13},{\omega}^{7},1,{\omega}^{3},{\omega}^{9},{\omega},{\omega}^{2},{\omega},{\omega}^{7},{\omega}^{6},{\omega}^{12},{\omega}^{11}).$

### v2

@@ III-B The case of NMDS ESGRS codes
    If $2\nmid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ over $\mathbb{F}_{q}$ with
    $\mathcal{A}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\}$
    and
    $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty).$
@@
    If $2\mid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ over $\mathbb{F}_{q}$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)$
    and
    $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1).$
@@
1) Since $\mathcal{S}$} contains a $k$-zero-sum subset, it follows from Lemma 3.2) that $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1$. From $2\nmid(n-k)$ and $3\leq k\leq n-2$, we get that $3\leq k\leq n-3$ and $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k-1}{2}\leq\frac{n-4}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell-1,2\ell+2]_{q}$. In this case, we immediately derive from (4) that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ has a generator matrix of the form
@@
Take
$\mathcal{A}$ $=\{(c_{1},\ldots,c_{n},0):~(c_{1},\ldots,c_{n})\in{\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1})\}$
$=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\}$
and $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)$. According to (1), we deduce that $\mathcal{A}$ and $\mathcal{B}$ have generator matrices of the forms
$G_{\mathcal{A}}=\begin{pmatrix}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a^{\ell}_{1}&\ldots&a^{\ell}_{n}&0\end{pmatrix}$
and
$G_{\mathcal{B}}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-1}&1\\ \end{pmatrix}$,
@@
$d(\mathcal{B}^{\perp})=d({\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)^{\perp})=\ell+1>\ell$
$\dim(\mathcal{A})=\ell+1>\ell,$
and
$d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$ $=n+1-(\ell+1)+2\ell+2$
$=n+\ell+2$
$>n+1,$
@@
$\left(\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\},{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ over $\mathbb{F}_{q}$. This completes the proof of 1).
@@
$d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$ $=n+1-(\ell+1)+1+2\ell+1$
$=n+\ell+2$
$>n+1.$
@@
$\left({\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ over $\mathbb{F}_{q}$ by Definition 3. This completes the proof of 2). ∎
@@
Note that we suppose $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$ for a more convenient representation in Theorems 9 and 10. Combining with Lemma 3, we can further deduce the explicit values of $\ell$ as listed in Table I, where the notation is the same with that used in Theorems 9 and 10.
TABLE I: $\ell$-ECPs $(\mathcal{A},\mathcal{B})$ of ESGRS codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$
@@
IV Unique decoding of ESGRS codes based on their ECPs
In this section, we present an explicit unique decoding algorithm for ESGRS codes based on their $\ell$-ECPs established in Section III and discuss its decoding complexities. We also present some concrete examples to illustrate our decoding algorithm.
IV-A An explicit unique decoding algorithm of ESGRS codes
In this subsection, we provide an explicit unique decoding algorithm for ESGRS codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on their $\ell$-ECPs. First of all, we recall the error-correcting mechanism of $\ell$-ECPs for a given $n,k]_{q}$ linear code $\mathcal{C}$} (see also [13, 7]). Let $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of $\mathcal{C}$} and let ${\bf y}={\bf c}+{\bf e}\in\mathbb{F}_{q}^{n}$ be a received vector with ${\bf c}\in{\mathcal{C}}$ and ${{\rm{wt}}}({\bf e})\leq\ell$. Note that based on conditions $\rm i)$$\rm iii)$} in Definition 3, Duursma and Kötter showed in [8, Theorem 1] that there exists a nonzero vector ${\bf a}\in\mathcal{A}$ such that
@@ IV-A An explicit unique decoding algorithm of ESGRS codes
has a unique solution, where $H_{{\mathcal{C}}}$ is a parity-check matrix of $\mathcal{C}$}. It then follows that ${\bf x}={\bf e}$ and $x_{i}=0$ for all $i\notin z({\bf a})$ if we write ${\bf x}=(x_{1},\ldots,x_{n})$. As a result, the $\ell$-ECP $(\mathcal{A},\mathcal{B})$ has uniquely decoded the received vector $\bf y$} to the codeword ${\bf c}={\bf y}-{\bf x}$.
Based on the above discussion, we conclude that to decode a linear code $\mathcal{C}$} using its $\ell$-ECP $(\mathcal{A},\mathcal{B})$, one should first determine explicit representations of the code pair $\mathcal{A}$ and $\mathcal{B}$, and then solve the system of equations (10) and (11). By combining Theorems 9 and 10, we can now present a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on its $\ell$-ECP as follows.
@@
If $\mathcal{S}$} is $k$-zero-sum free and $2\nmid(n-k)$, then an ESGRS code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ can be uniquely decoded by Algorithm 1 in $O(n^{3})$ time. Moreover, the decoding algorithms with the same time complexity for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ with respect to the other three cases discussed in Theorems 9 and 10 can be similarly derived by modifying Lines 8-14 of Algorithm 1 based on Table I accordingly.
@@
The unique decoding correctness is straightforward by combining Theorems 9 and 10 as well as the discussions above. By [30, Remark 2.2(1)], the algorithm takes $O((nt)^{3})$ time, with $t$ denoting the degree of the field extension of the finite field that contains the $\ell$-ECP $(\mathcal{A},\mathcal{B})$ over the base field of the ESGRS code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. By Theorems 9 and 10, we have $t=1$ and hence, the time complexity reduces to $O(n^{3})$. More details can be found in Algorithm 1. ∎
IV-B More discussions on the decoding complexity
In Theorem 11, we showed that the time complexity of Algorithm 1 is $O(n^{3})$ in general. Similar to [18], [35], [37], and [41], the extended Euclidean algorithm could also be a potential method to further deduce a unique decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ that operates in $O(qn)$ time. Since our algorithm complexity only depends on the code length $n+1$ and $n\leq q$, it is more efficient than those based on the extended Euclidean algorithm when $q>n^{2+\epsilon}$ for any $\epsilon>0$. On the other hand, the extended Euclidean algorithm starts from the parity-check matrix, but according to Lemma 2, the dual code of an ESGRS code may not be an ESGRS code of the same form. Therefore, even when the code length is linearly related to $q$, additional work must be done to precisely implement the extended Euclidean algorithm. In this sense, our decoding algorithm based on $\ell$-ECPs can achieve better performance than the extended Euclidean algorithm.
Moreover, focusing on the system of linear equations
which appears in Line 15 of Algorithm 1, we know from [8, Section VI.B] that there are two potential approaches to reduce the decoding complexity of Algorithm 1 to $O(n^{2})$ or even lower:
According to Theorems 9 and 10, the matrices $G_{\mathcal{A}}$ and $G_{\mathcal{B}}$ are closely related to the choices of $\mathcal{S}$} and $\bf v$}, rather than anything else. Note also that ${\mathrm{diag}}({\bf y})$ is a diagonal matrix only depending on the error vector $\bf e$} with ${\bf y}={\bf c}+{\bf e}$. Therefore, it is possible to design $\mathcal{S}$} and $\bf v$} or consider some special error patterns such that the above two approaches can be reached, leading to more efficient computations of solving the system of linear equations $G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}\cdot{\bf s}^{T}={\bf 0}.$
@@ 11    ${{\mathbf{v}}}^{\prime}\leftarrow(v_{1}(a_{1}-\gamma)^{k-1},\ldots,v_{n}(a_{n}-\gamma)^{k-1},1)$
12    ${{\mathbf{u}}}^{\prime}\leftarrow(u_{1}^{\prime},\ldots,u_{n+1}^{\prime})$ with $u_{i}^{\prime}$ given in Theorem 9.1)
@@ 26          ${\bf c}\leftarrow{\bf y}-{\bf x}_{0}$
Algorithm 1 An algorithm for uniquely decoding ${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)$ when $\mathcal{S}$} is $k$-zero-sum free and $2\nmid(n-k)$
IV-C Decoding examples of success and failure
@@
Let $q=17$, ${\mathcal{S}}=\{1,3,5,7,10,12,14,16\}\subseteq\mathbb{F}_{17}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1)\in(\mathbb{F}_{17}^{*})^{8}$. Consider the ESGRS code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ with generator matrix
@@
Suppose that the received vector is ${\bf y}=(4,6,1,14,5,7,12,15,\\ 2)\in\mathbb{F}_{17}^{9}$ and we have
@@
  • Note that $\mathcal{S}$} is $3$-zero-sum free. It then follows from Lemmas 2, 3.1), and 4 that the ESGRS code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $9,3,7]_{17}$ non-GRS MDS code with parity-check matrix
@@
    $G_{\mathcal{A}}=\begin{pmatrix}1&1&1&1&1&1&1&1&1\\ 16&1&6&7&15&12&10&11&0\\ 1&1&2&15&4&8&15&2&0\\ \end{pmatrix}$
    and
    $G_{\mathcal{B}}=\begin{pmatrix}5&14&13&3&12&13&5&3&3\\ 12&14&10&4&10&3&16&16&0\end{pmatrix}$.
@@
    As a result, we get the solution ${\bf x}_{0}=\{0,16,0,0,0,6,0,0,\\ 0\}$ and ${{\rm{wt}}}({\bf x}_{0})=2$.
  • Finally, according to Lines 24 and 25 of Algorithm 1, the received vector $\bf y$} is then uniquely decoded to the codeword
@@
Let $q=16$ and let $\omega$} be a primitive element of $\mathbb{F}_{16}$ such that ${\omega}^{4}+{\omega}+1=0$, ${\mathcal{S}}=\{1,{\omega},{\omega}^{2},{\omega}^{3},{\omega}^{4},{\omega}^{6},{\omega}^{7},{\omega}^{8},{\omega}^{9},{\omega}^{11},{\omega}^{12},{\omega}^{13},{\omega}^{14}\}\subseteq\mathbb{F}_{16}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1,1,1,1,1,1)\in(\mathbb{F}_{16}^{*})^{13}$. Consider the ESGRS code ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$ with generator matrix of the form (12).
$G_{7}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{3}&{\omega}^{4}&{\omega}^{6}&{\omega}^{7}&{\omega}^{8}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{13}&{\omega}^{14}&0\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{6}&{\omega}^{8}&{\omega}^{12}&{\omega}^{14}&{\omega}&{\omega}^{3}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{13}&0\\ 1&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&0\\ 1&{\omega}^{4}&{\omega}^{8}&{\omega}^{12}&{\omega}&{\omega}^{9}&{\omega}^{13}&{\omega}^{2}&{\omega}^{6}&{\omega}^{14}&{\omega}^{3}&{\omega}^{7}&{\omega}^{11}&0\\ 1&{\omega}^{5}&{\omega}^{10}&1&{\omega}^{5}&1&{\omega}^{5}&{\omega}^{10}&1&{\omega}^{10}&1&{\omega}^{5}&{\omega}^{10}&0\\ 1&{\omega}^{7}&{\omega}^{14}&{\omega}^{6}&{\omega}^{13}&{\omega}^{12}&{\omega}^{4}&{\omega}^{11}&{\omega}^{3}&{\omega}^{2}&{\omega}^{9}&{\omega}&{\omega}^{8}&1\\ \end{array}\right)$ (12)
@@
${\mathrm{diag}}({\bf y})={\mathrm{diag}}({\omega}^{12},{\omega},$ ${\omega}^{13},{\omega}^{7},1,{\omega}^{3},$
${\omega}^{9},{\omega},{\omega}^{2},{\omega},{\omega}^{7},{\omega}^{6},{\omega}^{12},{\omega}^{11}).$